### v1

# Rusty Linux: Advances in Rust for Linux Kernel Development

Abstract.
Context: The integration of Rust into kernel development is a transformative endeavor aimed at enhancing system security and reliability by leveraging Rust’s strong memory safety guarantees. Objective: We aim to find the current advances in using Rust in Kernel development to reduce the number of memory safety vulnerabilities in one of the most critical pieces of software that underpins all modern applications. Method: By analyzing a broad spectrum of studies, we identify the advantages Rust offers, highlight the challenges faced, and emphasise the need for community consensus on Rust’s adoption. Results: Our findings suggest that while the initial implementations of Rust in the kernel show promising results in terms of safety and stability, significant challenges remain. These challenges include achieving seamless interoperability with existing kernel components, maintaining performance, and ensuring adequate support and tooling for developers. Conclusions: This study underscores the need for continued research and practical implementation efforts to fully realize the benefits of Rust. By addressing these challenges, the integration of Rust could mark a significant step forward in the evolution of operating system development towards safer and more reliable systems.
Memory safety, Rust, Kernel, Operating System, Linux
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: General and reference Surveys and overviewsccs: Software and its engineering Language featuresccs: Computer systems organization Architectures
1. Introduction
The 1995 movie Hackers prescient predictions regarding the ease of breaking into computing systems in cyberspace have come to fruition. The White House Office of the National Cyber Director (ONCD) released a report calling for the technical community to proactively reduce the attack surface in cyberspace with a two-pronged approach (United States Gov, 2024). First, we need to address the root cause of many of the most heinous cyber attacks: memory-unsafe programming languages. Second, we need to establish better cybersecurity quality metrics so we can have a better understanding of the cybersecurity landscape.
In the ever-evolving landscape of software development, the reliability and security of computer systems stand as a paramount concern for all parties involved. Modern software is constructed by building ever more complex abstractions, one on top of the other. Thus, if we aim to have a secure system, we must start to peel back all the layers and tackle one of the fundamental abstractions in computer science: the programming language. Programming languages that provide and enforce memory safety eliminate whole classes of bugs, such as buffer overflows, dangling pointers, and memory leaks which have been implicated in a myriad of security vulnerabilities and system crashes.
Most operating system kernels are predominantly written in the C programming language with bits of assembly. C has been favored for its low-level capabilities and performance efficiency, crucial for kernel development. However, the inherent lack of memory safety in C has led to numerous security vulnerabilities, including buffer overflows and use-after-free errors, which have plagued operation system development for decades (MSRC Team, 2019). Addressing these vulnerabilities is paramount to enhancing the security and reliability of all operating systems.
Rust, a systems programming language developed by Mozilla, has garnered significant attention for its strong emphasis on memory safety without sacrificing performance. Rust’s ownership model and its compile-time checks effectively prevent common programming errors that lead to security vulnerabilities (Klabnik and Nichols, 2024). What makes Rust unique is the fact that it accomplishes these tasks without the use of a garbage collector, making it an ideal candidate for kernel development where safety and performance are critical.
The potential integration of Rust into kernel development represents a significant shift in the landscape of operating system development. This shift prompts a comprehensive evaluation of both the opportunities and challenges associated with using Rust as such a foundational component of modern computing. Therefore, we aim to synthesize existing research and practical experiences related to the use of Rust in the kernel, providing a detailed understanding of the current state of this emerging field.
This reflection paper presents a systematic literature review (SLR) focusing on strategies and methodologies for integrating Rust into one of the most fundamental areas that are typically dominated by unsafe languages, the operating system kernel. We aim to provide a comprehensive overview of existing research, identify gaps, and suggest future directions in this domain. Through a rigorous search process, we synthesized relevant studies and extracted key findings to offer insights into effective approaches for ensuring memory safety when working closely with hardware.
Figure 1. Process Diagram for Paper selection
2. Research Methodology
For our research methodology, we followed the Kitchenham and Charters (Kitchenham, 2004) methodology to conduct SLR in software engineering. We divided our review into three discrete phases, planning the review, conducting the review, and reporting the review results. The following sections detail our review process and are diagrammed in figure 1.
2.1. Planning
To begin our study, we must first confirm the need for a SLR. The recent report released by ONCD (United States Gov, 2024) has conveniently done this job for us by compiling a report detailing the need for research in the domain of memory safety. While the ONCD report detailed a two-pronged approach, for this reflection paper, we will be focusing on a memory-safe programming language, specifically Rust. While there are many modern memory-safe programming languages available for software developers to use, Rust is one of the few languages that is feasible to use when developing operating system kernels due to its lack of runtime and garbage collector (Ojeda, 2021). The Linux kernel is currently in the early stages of adding real support for Rust (The kernel development community, [n. d.]) thus, we will focus on Rust as a primary candidate to replace the aging C programming language.
2.2. Research Questions
We have defined the following research questions:
  • RQ1: What are the existing approaches for implementing operating system kernels in Rust?
    Although Rust in the Linux kernel is recent, we want to see what approaches researchers have explored in any operating system. Experiential and research kernels can provide new ideas that can be integrated into the Linux kernel.
  • RQ2: What are the performance implications of using Rust for operating system kernel development?
    The C programming language has the advantage of 40 years of improvements and lessons learned regarding how to write fast code. Rust, being a new language, still has many unexplored areas regarding how to write fast code, especially in kernel space. Therefore, we want to explore the performance implications of Rust for kernel development in terms of throughput, latency, and resource utilization.
  • RQ3: What are the major challenges and limitations when developing operating system kernels in Rust?
    Kernel space presents unique challenges for the software engineer using Rust. Rust was not designed to implement kernels and is more restrictive than C. Thus, we want to identify if there are any stumbling blocks that have already been found when using Rust in kernel development.
  • RQ4: What are the lessons learned when developing operating systems kernels in Rust?
    Besides technical challenges and limitations, developers’ perceptions are very important in software development. Therefore, through this research question, we want to find developers’ reflections on what went well and what didn’t go well while developing kernels in Rust.
2.3. Data Collection
2.3.1. Search Strategy.
We employed a robust multi-step search strategy across four databases: ACM Digital Library, IEEE Xplore, Scopus, and Springer Link to find all the current research. We leveraged the advanced search features of all three databases to search the titles and abstracts for keywords using boolean search operators.
2.3.2. Search criteria.
We searched all four databases with the keywords listed in table 1 between January 1, 2019 and April 1, 2024. We choose to look at only the previous 5 years of research in order to capture the bleeding-edge research that is currently being done. While both ACM and IEEE had mutually exclusive results, Scopus had 44 duplicates that needed to be removed and Springer Link didn’t have any new papers that were not also in ACM or IEEE. We then used Google Scholar to do forward and backward snowballing.
Database Query Results
ACM (Abstract:(”operating system” kernel linux OS) AND Abstract:(rust) )OR (Title:(”operating system” OS kernel linux) AND Title:(rust)) 30
IEEE (”All Metadata”:”operating system” AND ”All Metadata”:rust) OR (”All Metadata”:kernel AND ”All Metadata”:rust) OR (”All Metadata”:linux AND ”All Metadata”:rust) 39
Scopus TITLE-ABS-KEY ( ( ”operating system” OR kernel OR linux ) AND rust ) AND PUBYEAR ¿ 2018 AND PUBYEAR ¡ 2025 AND ( LIMIT-TO ( SUBJAREA , ”COMP” ) ) AND ( LIMIT-TO ( LANGUAGE , ”English” ) ) 87
Springer Link Operating System Rust Kernel (Conference Paper, Article, Research article) Subdiscipline: Software engineering/programming and operating systems 29
Table 1. Search Queries used for each Database.
2.3.3. Inclusion and exclusion criteria.
Approach Papers Operating System in Rust
Monolithic (The kernel development community, [n. d.]), (Li et al., 2019), (Miller et al., 2021), (Oikawa, 2023) Linux kernel v6.1+
Micro-kernel (Chen et al., 2023),(Liang et al., 2021), (Liu et al., 2024), (Narayanan et al., 2020), (Narayanan et al., 2019) Atmosphere, Redox, Redleaf
Embedded (Culic et al., 2022), (Vishnunaryan and Banda, 2022) Tock, Hubris, Drone, Bern, HarSaRK
Unikernel (Lankes et al., 2019), (Boos et al., 2020), (Ijaz et al., 2023), (Sung et al., 2020) RustyHermit, Theseus
Exokernel (Li and Sato, 2024) W-Kernel
Table 2. Approaches and Methodologies for Rust in the Kernel
Following the guidelines outlined by Kitchenham and Charters (Kitchenham, 2004) we set the inclusion and exclusion criteria based on our research questions outlined in section 2.2. We only considered papers that are written in English and published in conferences, journals, and workshops. The published papers should describe using the Rust programming language for either developing a new kernel, extending an existing kernel, or authoring drivers. We included any type of kernel architecture, including monolithic kernel, microkernel, or unikernel in both the embedded and non-embedded space, as long as the paper is using the Rust programming language in some way. Papers that describe solutions that reside 100% in user space or papers that relied too heavily on a component written in C were excluded. Early efforts to use Rust in kernel development were impacted by changes and updates in both the Rust language and compiler and relied on unstable features, which were thus excluded from this review.
2.4. Data Analysis
Once we had our initial collection of papers from the databases, we merged all duplicate papers into one record and then started our first pass, which consisted of reading the title and abstract and either marking the paper as include or exclude. Once we completed the initial screening, we were left with 26 papers that needed an in-depth reading and analysis. We found an additional 6 papers during forward and backward snowballing that were also reviewed, bringing the total number of papers to 32. The first author manually reviewed all 32 papers, built the first dataset, and conducted the analysis. The second author then verified all the data and results. Since the results were straightforward and came from a comparatively small number of papers, there were no discrepancies or missing results to solve.
3. Results
In this section, we present our findings. We address each research question from RQ1 - RQ4.
3.1. RQ1: Existing Approaches
The integration of Rust into operating system kernel development has garnered significant attention due to Rust’s promise of memory safety and concurrency guarantees. Several approaches and methodologies have emerged to leverage Rust for kernel development, each addressing different aspects of kernel functionality and integration. We found several approaches that researchers are taking to integrate Rust into kernel development. The first approach is a greenfield approach where the operating system architecture is redesigned and built from the ground up in order to fully leverage the Rust programming language (Culic et al., 2022; Boos et al., 2020). The second approach is the incremental integration of Rust into existing kernels, such as the Linux kernel (The kernel development community, [n. d.]; Miller et al., 2021; Oikawa, 2023). This methodology involves gradually rewriting or augmenting specific components of the kernel with Rust while maintaining compatibility with the existing C codebase. Finally, researches are porting existing C kernels over to Rust and then evaluating the differences. Table 2 summarizes our findings regarding what existing approaches researchers are using to integrate rust into the kernel.
For the first approach, Boos et al. (Boos et al., 2020) created an experimental operating system named Theseus that operates in a single address space and single privilege level and uses properties of the Rust programming language to realize isolation instead of relying on hardware. This novel approach pushes some of the roles that a traditional operating system would take onto the compiler. The design of Theseus uses a novel cell-based structure where ownership of memory and resources is enforced by the compiler, thus avoiding these checks during runtime to get improved performance. In addition to memory safety, Theseus has a goal of addressing the state spill problem. State spill happens when a single service in an operating system can harbor a state change induced by interacting with other services. That state change can eventually cause a system crash or system instability at some point in the future, making it very difficult to track down the problem due to the separation between when the problem occurred and when it was observed. Theseus OS is an operating system that was designed to match the Rust language instead of the more traditional route of matching the hardware. Li and Sato (Li and Sato, 2024) explored using Rust to write an exokernel named W-Kernel. The authors proposed a novel architecture for an OS by embedding a WebAssembly (WASM) runtime into the kernel that can run programs written in any language that can compile down to WASM.
For the second approach, researchers take a more traditional micro-kernel implementation (Chen et al., 2023; Liang et al., 2021). By pushing most kernel functionality to user space, the authors limited the surface area that needs to be proved. Chen X et al. (Chen et al., 2023) took their implementation one step further by formally verifying their new micro-kernel named Atmosphere by using both the liner type system of Rust in combination with an SMT solver. They were able to get a 7.5:1 proof-to-code ratio, which is higher than other formerly verified micro-kernels SeL4 and CeriKOS both have proof code ratios of 19:1 and 20:1, respectively. While this approach was novel drivers running in userspace are still not formally verified and don’t have to be written in Rust.
Finally, researchers have taken existing Unikernels written in C and ported them to Rust. Unikernels are designed to do one thing and one thing only, thus limiting the complexity that a general-purpose operating system has to deal with (Madhavapeddy and Scott, 2014). This reduced complexity allowed Lankes et al. (Lankes et al., 2019) to take HermitCore and rewrite it in Rust. The new Unikernel named RustyHermit consists of only 3.27% unsafe Rust, with the rest of the code base consisting of safe Rust, dramatically decreasing the surface area where memory errors could originate.
3.2. RQ2: Performance Implications
No. Implication Studies that Reported the challenge
1 Performance (Gonzalez et al., 2023), (Li et al., 2024), (Ma et al., 2023)
2 Throughput (Gonzalez et al., 2023)
3 Latency (Culic et al., 2022)
Table 3. Performance Implications of Rust in the Kernel
Comparing performance, throughput and latency between different systems with different architectures is very difficult. For example, several of the operating systems written in Rust are a complete rethinking of how an operating system is designed (Boos et al., 2020) while others run entirely in kernel space (Lankes et al., 2019). Additionally, direct comparisons between a Monolithic kernel, Micro-kernel, and Unikernel, or comparisons between a Real Time Operation System (RTOS) and a General Purpose Operating System, are not directly meaningful due to the vastly different goals of each system and the overhead that is imposed by the hardware (Arnott, 2012). Therefore, we will focus more on the overall efforts that are specific to the Rust programming language and the challenges presented. We summarize our findings in table 3.
Culic et al. (Culic et al., 2022) looked at latency issues in Tock. Tock is a new operating system written in Rust that is designed to run on embedded systems but does not provide real-time capabilities. The authors attempted to add real-time capabilities by integrating eBPF into the Tock kernel to improve the interrupt handler’s response time. The authors found that early work (still in the prototype stage) lowers the response times of the system and interrupts response times 3x.
Gonzalez et al. (Gonzalez et al., 2023) explored using the Rust for Linux Project to implement a native UDP driver in Rust in order to explore the performance of the Rust programming language. The authors were able to get a basic driver working with performance only slightly slower than C using the Rust for Linux (RFL) project. The RFL project is still too immature to get a full driver up and running, but it is at a stage where researchers can start experimenting with different approaches.
Li et al. (Li et al., 2024) explored the feasibility of using Rust in kernel space. The authors took an existing component, the Out of Memory (OOM), and implemented a replacement using the Rust programming language. The non-encapsulated interface Rust component which was almost identical to the original C component, only introduced a 0.7% overhead. The encapsulated Rust component, on the other hand, added a 3% performance overhead.
3.3. RQ3: Challenges and Limitations
No. Challenge Description
1 Binary Size (Ayers et al., 2022) Deeply ingrained monomorphization which increased the size of the rust binaries Compiler optimizations are not as mature as some C based compilers, thus increasing binary size Hidden data structures and data Sub-optimal compiler generated support code
2 Missing Features (Burtsev et al., 2021) Supporting trait bounds on functions and closures with any number of arguments Expose type information in procedural macros Support a collision-free, unique type identifier Support typed assembly language for Rust Support trusted build environments Provide software-only stack guard with extensible probing interface Develop zero-copy serialization of “plain-old” data structures
3 Soundness (Klimt et al., 2023) Unsynchronized Global State - any use of mutable statics is unsafe C-Style Abstractions - Use rust style abstractions to properly encapsulate internal unsafe usage of raw pointers Aliasing Mutable References - Giving out raw pointers to memory that is also referenced mutably Re-implementing Memory Access - instead of accessing specific memory regions in assembly, creating and using references to the whole region is preferable
4 Panics (Ma et al., 2023),(Burtsev et al., 2021) Support extendable, no_std unwind library Stack unwinding in embedded environments
5 C Interop (Miller et al., 2019), (Li et al., 2021) Kernel interfaces, while designed for extensibility, are not designed for type safety Hybrid Code Flow. The Rust compiler can not track ownership when switching between modules written in C and Rust
Table 4. Challenges Unique to the Rust Programming Language
The development of operating system kernels in Rust introduces several unique challenges and limitations compared to more traditional languages like C. These challenges arise from Rust’s strict safety guarantees, its relatively recent adoption in systems programming, and the inherent complexities of kernel development (Ojeda, 2021). This section discusses these challenges, limitations, and the lessons learned from various projects and research efforts. We summarize our findings in table 4.
A primary concern for programming languages that are used for operating system development is the size of the binary especially in an embedded environment, we must be careful to keep the size of the Rust component as close as we can to the C component so we can still run on the same hardware. Li et al. (Li et al., 2024) found that there was only a 0.06% size increase when compared to the original C implementation. Ayers et al. (Ayers et al., 2022) focused on reducing the size of binaries produced by the rust compiler while working with Tock OS. They were able to identify several causes of binary growth that are specific to the Rust programming language and have the following 5 recommendations when using Rust in a size-constrained environment:
  • Minimize Length + Instantiations of Generic Code
  • Use Trait Objects Sparingly
  • Don’t Panic
  • Carefully Use Compiler Generated Support Code
  • Don’t use static mut
Burtsev et al. (Burtsev et al., 2021) explored what is missing in the Rust programming language to help solve the isolation problem in an operating system. Currently, Rust lacks the ability to express isolation in the heap without external support. For example, the RedLeaf experiential operating system, which was written in Rust, relied on a complex interface definition language (IDL) to enforce isolated heaps. The paper enumerates several properties of the Rust Language that could help with isolation with regards to operating system development that are detailed in table 4. The authors argue that with these changes or inclusions to the Rust programming language developing operating system kernels would be much easier and safer.
3.4. RQ4: Lessons Learned
Klimt et al. (Klimt et al., 2023) details the lessons learned and challenges when implementing Theseus. First, they found that it is impossible to write a complete operating system in 100% safe Rust. For example, when writing a memory management system, raw pointers must be used to modify the hardware. Despite the limitations imposed by the hardware, the authors describe how Theseus leveraged intralingual design to maximize the compiler’s role in enforcing correctness. By leveraging Rust’s type system and borrow checker memory safety, correct ownership transfer can be achieved at a higher level than what could be done in C. The authors also detailed some of the limitations of intralingual design, such as not being as expressive as many other formal verification techniques due to the limited invariance that the type system can enforce. One of the most important lessons learned was the insight that a linear type system itself cannot guarantee the uniqueness of the resource represented, such as when a memory resource may overlap. The authors introduce the idea of using a hybrid approach of verification where they leverage both the linear type system and an SMT solver. The authors also explore the bootstrapping problem with rust systems. In operating systems written in C, the kernel provides the ownership root to applications running on top. Finding a new ownership root is an open research question with regard to using Rust for operating system development.
4. Threats to Validity
Internal Threats. Potential internal threats include:
  • Quality of Studies: The methodological rigor of the included studies varies, and some may suffer from design flaws or biases that were not adequately controlled for, thus affecting the reliability of their findings.
  • Selection Criteria: The inclusion and exclusion criteria for selecting relevant studies might inadvertently bias the review towards certain types of research, such as those reporting successful integration while underrepresenting studies detailing challenges and failures.
External Threats. Threats to external validity include:
  • Context-Specific Findings: Many studies focus on specific kernel modules or use-cases, which may not be representative of the broader kernel environment. The success of Rust in isolated components does not necessarily translate to the entire kernel.
  • Temporal Changes: The rapidly evolving nature of both the Rust programming language and the Linux kernel means that findings from older studies may no longer be applicable, as improvements and changes in both domains can alter the landscape significantly.
  • Definition Ambiguity: The term “Kernel development in Rust” encompasses a wide range of activities, from minor module development to complete subsystem rewrites. Variations in how researchers and developers interpret this integration could lead to inconsistent findings.
Conclusion Threats. Potential conclusion threats include:
  • Heterogeneity of Studies: The diverse methodologies, metrics, and contexts of the included studies can lead to challenges in synthesizing findings and drawing unified conclusions. For example, it is very difficult to directly compare a micro-kernel, Unikernel, and hybrid Kernel.
  • Reviewer Bias: Personal biases of the reviewers in interpreting data and making judgments about study quality and relevance may skew the results
We made efforts to minimize this through predefined criteria and multiple reviewers, but some degree of subjectivity is inevitable. In addressing these threats, we employed rigorous methods for study selection, data extraction, and analysis, and we remain transparent about the limitations. Despite these threats, our results provide a valuable synthesis of the current state of research on the use of Rust in kernel development, highlighting both its potential benefits and the challenges that need to be addressed.
5. Discussion & Conclusion
The integration of Rust into kernel space represents a significant evolution in operating system development, promising to address long-standing issues related to memory safety and system reliability. We synthesized the current state of research and practice concerning the use of Rust in multiple types of kernels, including the venerable Linux kernel, highlighting both the progress made and the challenges that remain.
Our review indicates that Rust’s strong guarantees of memory safety, enabled by its ownership model and strict compile-time checks, offer a compelling advantage over traditional C-based kernel development. These features have the potential to reduce common vulnerabilities such as buffer overflows and use-after-free errors, which are prevalent in C and have historically led to critical security exploits. Several case studies and prototype implementations have demonstrated that Rust can be successfully integrated into the kernel, providing safer interfaces and reducing the incidence of memory-related bugs without incurring significant performance penalties.
However, the adoption of Rust for kernel development is not without its challenges. Truly, the biggest challenge is existing codebases and mountains of legacy code. With extensive use of C and reliance on specific C idioms and low-level programming techniques, researchers have substantial integration hurdles to overcome. Any efforts to rewrite substantial portions of an existing kernel such as Linux in Rust, are constrained by the need for interoperability with existing C code and the necessity to maintain the kernel’s performance characteristics. Furthermore, the operating system development community must reach a consensus on Rust’s role and ensure that sufficient tooling, documentation, and support are available for developers.
In conclusion, while Rust’s incorporation into the kernel space is still in its nascent stages, the initial results are promising. The potential for enhanced security and stability aligns with the long-term goals of kernel development, and continued research, coupled with practical implementation efforts, will be crucial in realizing these benefits. Future work should focus on addressing integration challenges, refining interoperability mechanisms, and expanding the body of empirical evidence on Rust’s impact within the kernel environment. As the community navigates these challenges, the evolution of Rust in the kernel may well mark a transformative period in the pursuit of safer, more reliable operating systems.